# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v2

@@ 3.2 Contraction and mixing
$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}\qed$
Part II: Extension to constrained distributions.
We now prove a technical lemma that extends the contraction results to constrained distributions. This lemma guarantees the existence of weak solutions to two stochastic processes that describe the evolution of distributions involved in Line 2 and 3 in $\mathsf{In\text{-}and\text{-}Out}$, in addition to lower-semicontinuity of $f$-divergence. We shall prove it for any measure that is absolutely continuous with respect to $\pi^{X}$, since this imposes no additional technical hurdles.
Lemma 22.
Let $\nu$ be a measure, absolutely continuous with respect to the uniform measure $\pi^{X}$. The forward and backward heat flow equations given by
$\partial_{t}\mu_{t}$ $=\frac{1}{2}\Delta\mu_{t}\,,$
$\partial_{t}\mu_{t}^{\leftarrow}$ $=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\nu P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{with }\mu_{0}^{\leftarrow}=\mu_{h}\,,$
admit solutions on $(0,h]$, and the weak limit $\lim_{t\to h}\mu_{t}^{\leftarrow}=\mu_{h}^{\leftarrow}$ exists for any initial measure $\mu_{0}$ with bounded support. Moreover, for any $f$-divergence with $f$ lower semi-continuous,
$D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu_{t})\,.$
Proof.
The existence of weak solutions for the forward equation is well-known, since $\mu_{0}$ can be weakly approximated by measures with continuous density, for which the heat equation admits a unique solution for all time. In particular, the weak solution is $C^{\infty}$ for $t>0$.
The reverse SDE is more subtle, since $\nabla\log\nu P_{t}$ will in general cease to be Lipschitz as $t\to 0$. On the other hand, for any $h>0$, we can write explicitly
$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$
If one considers the system started at $\tilde{\mu}_{0}=\mu_{\epsilon}=\nu P_{\epsilon}$ and solve the forward-backward Fokker-Planck equations on times $[0,h-\epsilon]$, then $\tilde{\mu}_{h-\epsilon}=\mu_{h}=\mu_{0}^{\leftarrow}=\tilde{\mu}_{0}^{\leftarrow}$ and
$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in §D.
We note that $\tilde{\mu}_{h}^{\leftarrow}$ defined below is obviously integrable with respect to $x$, since this is a probability density function of $\mu_{h}Q$, where $Q$ is the backward kernel:
$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
Hence, $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined over the support of $\nu$ almost surely.
We now show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$. Note that $\nu P_{\epsilon}\to\nu$, as $\nu$ has a bounded support. Also, the denominator is independent of $\epsilon$ due to
$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}\bigl{(}0,(h-\epsilon)I\bigr{)}*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
Hence, for $\epsilon\leq d^{-1}$,
$\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}$ $\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}$
$\lesssim\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
Since the last bound is integrable with respect to $\mu_{h}$, the dominated convergence theorem implies
$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
Thus, the pointwise convergence follows. Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backward step of our algorithm. In particular, this corresponds to first sampling $x\sim\mu_{h}$, then $y\sim Q_{h}^{\nu,h}(\cdot|x)$, which is precisely the law of $\mu_{h}^{\leftarrow}$ given by ($\mathsf{FP\text{-}BH}$).
As for the second statement, it follows from Scheffé’s lemma [Bil95, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence. It follows from lower semicontinuity of $D_{f}$ [AFP00, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$. ∎
In the sequel, we will only consider $\nu=\pi^{X}$. Since the Rényi divergence is a continuous function of the $\chi^{q}$ divergence (see Definition 6), which itself is an $f$-divergence, it enjoys the same lower-semicontinuity properties. Using this lower-semicontinuity together with the decay results in Theorem 21, we can easily derive the contraction results of $\mathsf{In\text{-}and\text{-}Out}$ in $\mathcal{R}_{q}$ and $\chi^{q}$ for any $q\geq 1$. We remark that this result does not require convexity of $\mathcal{K}$.
Theorem 23.
Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$. Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$. Then, for any $q\geq 1$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{2k/q}}\,.$
For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$,
$\chi^{2}(\mu^{X}_{k}\mathbin{\|}\pi^{X})\leq\frac{\chi^{2}(\mu^{X}_{0}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{PI}})^{2k}}\,.$
Furthermore, for any $q\geq 2$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
Proof.
Let us set $\mu_{0}=\mu_{0}^{X}$ and $\pi_{0}=\pi^{X}$. Then, $\mu_{h}=\mu_{0}^{\leftarrow}=\mu_{1}^{Y}$, $\pi_{h}=\pi_{0}^{\leftarrow}=\pi^{Y}$, and $\mu_{h}^{\leftarrow}=\mu_{1}^{X}$, $\pi_{h}^{\leftarrow}=\pi^{X}$. For small $\epsilon>0$, as $\mu_{\epsilon}=(\mu_{0}^{X})_{\epsilon}=\mu_{0}^{X}*\mathcal{N}(0,\epsilon I_{d})$ is $C^{\infty}$-smooth, we can now invoke the decay results with step size $h-\epsilon$ in Theorem 21. Thus, for contraction constants $C_{\epsilon}=(1+\frac{h-\epsilon}{C_{\mathsf{LSI}}+\epsilon})^{-2/q}$ and $C_{\epsilon}=(1+\frac{h-\epsilon}{C_{\mathsf{PI}}+\epsilon})^{-2}$ respectively when $\Phi=\mathcal{R}_{q}$ and $\Phi=\chi^{2}$,
$\Phi(\mu_{h-\epsilon}^{\leftarrow}\mathbin{\|}\pi_{\epsilon})\leq C_{\epsilon}\cdot\Phi(\mu_{\epsilon}\mathbin{\|}\pi_{\epsilon})\leq C_{\epsilon}\cdot\Phi(\mu_{0}\mathbin{\|}\pi_{0})\,,$
where we used the data-processing inequality for the last inequality. By the second result of Lemma 22, sending $\epsilon\to 0$ leads to
$\Phi(\mu_{1}^{X}\mathbin{\|}\pi^{X})=\Phi(\mu_{h}^{\leftarrow}\mathbin{\|}\pi_{0})\leq C\cdot\Phi(\mu_{0}\mathbin{\|}\pi_{0})=C\cdot\Phi(\mu_{0}^{X}\mathbin{\|}\pi^{X})\,.$
Repeating this argument $k$ times completes the proof. ∎
@@ 3.3 Failure probability and wasted steps
We begin by defining a suitable version of local conductance [KLS97].
@@
Naïve analysis for expected number of trials.
Starting from $\pi^{X}$, when we just naïvely sample from $\pi^{Y|X}(\cdot|x)$ for all $x$ without imposing any failure condition, the expected number of trials for one iteration is that for the probability density $p_{x}$ of $\mathcal{N}(x,hI_{d})$,
@@
Refined analysis under a failure condition.
Going forward, we assume an $M$-warm start as in previous work for uniform sampling algorithms. By induction we have $\frac{\mathrm{d}\mu^{X}_{i}}{\mathrm{d}\pi^{X}}\leq M$ for all $i$.
@@
ensures that $\mathsf{F}$ vanishes at $s=\infty$. Hence, bounding the first term in (3.4) by 00 results in
@@
Now we choose a suitable threshold $N$ for bounding the failure probability. Following (3.3) in the proof, one can notice that for $y\in\mathcal{K}^{c}_{\delta}$, $\delta=\Omega(1/d)$, and $h=\Theta(d^{-2})$,
@@
Thus, the expected number of trials from $\mathcal{K}^{c}_{\delta}$ for the rejection sampling in Line 3 is $\ell(y)^{-1}\geq\exp(\Omega(t^{2}))$. Intuitively, one can ignore whatever happens in $\mathcal{K}^{c}_{\delta}$, since $\mathcal{K}_{\delta}$ takes up most of measure of $\pi^{Y}$. As the number of trials from $\mathcal{K}_{\delta}^{c}$ is at least $\exp(\Omega(t^{2}))$ in expectation, the most straightforward way to ignore algorithmic behaviors from $\mathcal{K}_{\delta}^{c}$ is simply to set the threshold to $N=\widetilde{\mathcal{O}}(\exp(t^{2}))$. Even though the threshold is $N$, the expected number of trials is much lower.
@@
For $\mu_{h}:=\mu*\mathcal{N}(0,hI_{d})$, the failure probability is $\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]$. Since $\mathrm{d}\mu/\mathrm{d}\pi^{X}\leq M$ implies $\mathrm{d}\mu_{h}/\mathrm{d}(\pi^{X})_{h}=\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$, it follows that
@@
where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}\bigl{(}(1+\delta)\mathcal{K}\bigr{)}=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$. Taking $c=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$. Therefore,
$\mathbb{E}_{\mu_{h}}[(1-\ell(\cdot))^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell(\cdot))^{N}]\leq\frac{\eta}{m}\,.$
@@
$\leq e^{t}+N\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\bigl{(}\log^{4}\frac{mM}{\eta}\bigr{)}\,.$
Therefore, the expected number of trials per step is $\mathcal{O}(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
@@ 3.4 Putting it together
We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in §1.2.
@@
For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and $N=\widetilde{\mathcal{O}}(\frac{mM}{\eta})$. For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations. With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
@@
We just put together Lemma 14 and Theorem 23. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (7) for $q\geq 2$ in Theorem 23. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$ iterations. Since each iteration has $\eta/m$-failure probability by Lemma 14, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations. Lastly, each iteration requires $\widetilde{\mathcal{O}}(M\log^{4}\nicefrac{{1}}{{\eta\varepsilon}})$ membership queries in expectation by Lemma 14. Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}})\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$ expected number of membership queries over $m$ iterations. Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
@@
Proof of Theorem 5..
By the Rényi-decay under (8) in Theorem 23, $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\varepsilon$-distance to $\pi^{X}$ after $\mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\mathcal{R}_{q}(\mu^{X}_{1}\mathbin{\|}\pi^{X})}{\varepsilon}\bigr{)}$ iterations for $q\geq 1$.
For $q\geq 2$, we use the decay result under (7). In this case, $\mathsf{In\text{-}and\text{-}Out}$ decays under two different rates depending on the value of $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$. It first needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X}))$ iterations until $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$ reaches $1$. Then, $\mathsf{In\text{-}and\text{-}Out}$ additionally needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{1}{\varepsilon})$ iterations, and thus it needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})+\log\frac{1}{\varepsilon}\bigr{)})$ iterations in total. By substituting $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$, we complete the proof. ∎

### v3

@@ 3.2 Contraction and mixing
$1-\exp\bigl(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr)\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
This completes the proof. ∎
Using the contraction result above, we conclude the mixing time of $\mathsf{In\text{-}and\text{-}Out}$.
Proof of Theorem 12.
Note that one iteration of $\mathsf{In\text{-}and\text{-}Out}$ corresponds to the composition of two transition kernels $P_{h}$ and $Q_{h}$, where $P_{h}(x,\cdot)=\mathcal{N}(x,hI_{d})$. The forward step ($\mathsf{FH}$) (i.e., transition kernel $P_{h}$) convolves the law $\mu_{k}^{X}$ of the $k$-th iterate with Gaussian $\gamma_{h}$, so the forward step invokes the contraction with rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$ by Lemma 19. Using the DPI (Lemma 9) to address the effect of $Q_{h}$, we obtain that
$\mathcal{R}_{q}(\mu_{k+1}^{X}\mathbin{\|}\pi^{X})=\mathcal{R}_{q}(\mu_{k+1}^{Y}Q_{h}\mathbin{\|}\pi^{Y}Q_{h})\underset{\text{DPI}}{\leq}\mathcal{R}_{q}(\mu_{k+1}^{Y}\mathbin{\|}\pi^{Y})\leq\frac{\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}}(\pi^{X}))^{1/q}}\,.$
Induction on $k$ proves the first claim. The other results can be proven in the similar fashion. ∎
@@ 3.3 Failure probability and wasted steps
We begin by defining a Gaussian version of local conductance [KLS97].
@@
Naïve analysis for expected number of trials.  Starting from $\pi^{X}$, when we just naïvely sample from $\pi^{Y|X}(\cdot|x)$ for all $x$ without imposing any failure condition, the expected number of trials for one iteration is that for the probability density $p_{x}$ of $\mathcal{N}(x,hI_{d})$,
@@
Refined analysis under a failure condition.  Going forward, we assume an $M$-warm start as in previous work for uniform sampling algorithms. By induction we have $\frac{\mathrm{d}\mu^{X}_{i}}{\mathrm{d}\pi^{X}}\leq M$ for all $i$.
@@
ensures that $\mathsf{F}$ vanishes at $s=\infty$. Hence, bounding the first term in (3.3) by 0 results in
@@
Now we choose a suitable threshold $N$ for bounding the failure probability. Following (3.2) in the proof, one can notice that for $y\in\mathcal{K}^{c}_{\delta}$, $\delta=\Omega(t/d)$, and $h=\Theta(d^{-2})$,
@@
Thus, the expected number of trials from $\mathcal{K}^{c}_{\delta}$ for the rejection sampling in Line 3 is $\ell(y)^{-1}\geq\exp(\Omega(t^{2}))$. Intuitively, one can ignore whatever happens in $\mathcal{K}^{c}_{\delta}$, since $\mathcal{K}_{\delta}$ takes up most of measure of $\pi^{Y}$. As the number of trials from $\mathcal{K}_{\delta}^{c}$ is at least $\exp(\Omega(t^{2}))$ in expectation, the most straightforward way to ignore algorithmic behaviors from $\mathcal{K}_{\delta}^{c}$ is simply to set the threshold to $N=\widetilde{O}(\exp(t^{2}))$. Even though the threshold is $N$, the expected number of trials is much lower.
@@
For $\mu_{h}:=\mu*\mathcal{N}(0,hI_{d})$, the failure probability is $\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]$. Since $\mathrm{d}\mu/\mathrm{d}\pi^{X}\leq M$ implies $\mathrm{d}\mu_{h}/\mathrm{d}(\pi^{X})_{h}=\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$ (as easily checked by the definition of convolution), it follows that
@@
where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}((1+\delta)\,\mathcal{K})=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$. Taking $c=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z\,(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$. Therefore,
$\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell)^{N}]\leq\frac{\eta}{m}\,.$
@@
$\leq e^{t}+N\exp\bigl(-\frac{t^{2}}{2c}+t\bigr)\leq\log^{3}Z+3\log^{4}Z=O\bigl(\log^{4}\frac{mM}{\eta}\bigr)\,.$
Therefore, the expected number of trials per step is $O(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
@@ 3.4 Putting it together
We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as the $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in §1.2.
@@
For any given $\eta,\varepsilon\in(0,1)$, $q>1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, step size $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and threshold $N=\widetilde{O}(\frac{mM}{\eta})$. For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations. With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{O}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
@@
We just put together Lemma 13 and Theorem 12. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (7) for $q\geq 2$ in Theorem 12. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=O(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 17), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ if it iterates at least
$m\gtrsim qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log d\log\frac{M}{\eta\varepsilon}\log\frac{mM}{\eta}\quad\text{times.}$
Note that $x\geq A\log Bx$ for $A,B\geq 1$ is satisfied when $x\gtrsim A\log AB=\widetilde{O}(A\log B)$. Using this, we set
$m=\widetilde{O}\bigl(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log\frac{M}{{\eta\varepsilon}}\log\frac{M}{\eta}\bigr)=\widetilde{O}\bigl(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{{\eta\varepsilon}}\bigr)\,.$
Since each iteration has $\eta/m$-failure probability by Lemma 13, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations. Lastly, each iteration requires $\widetilde{O}(M\log^{4}\nicefrac{{1}}{{\eta\varepsilon}})$ membership queries in expectation by Lemma 13. Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$ expected number of membership queries over $m$ iterations. Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=O(\log d)$. ∎
@@
Proof of Theorem 5.
By the Rényi-decay under (8) in Theorem 12, $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\varepsilon$-distance to $\pi^{X}$ after $O\bigl(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\mathcal{R}_{q}(\mu^{X}_{1}\mathbin{\|}\pi^{X})}{\varepsilon}\bigr)$ iterations for $q\geq 1$.
For $q\geq 2$, we use the decay result under (7). In this case, $\mathsf{In\text{-}and\text{-}Out}$ decays under two different rates depending on the value of $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$. It first needs $O(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\,\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X}))$ iterations until $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$ reaches $1$. Then, $\mathsf{In\text{-}and\text{-}Out}$ additionally needs $O(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{1}{\varepsilon})$ iterations, and thus it needs $O(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\bigl(\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})+\log\frac{1}{\varepsilon}\bigr))$ iterations in total. By substituting $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$, we complete the proof. ∎